OpenStudy (anonymous):

write a mathematical expression for each phrase, and simplify the expression. 1) four-fifths of the sum of -20 and 5

OpenStudy (agreene):

\[\frac{4}{5}(-20+5)=\frac{4}{5}(-15)=-12\]

OpenStudy (anonymous):

4/5(-20 + 5)

OpenStudy (anonymous):

theres no way to simplify it right?

OpenStudy (anonymous):

oh no nevermind!!

OpenStudy (agreene):

it doesnt get simpler than what I wrote (-12)

OpenStudy (anonymous):

what about this one. 12 less than the product of -3 and -6 ? is it -3+-6-12?

OpenStudy (anonymous):

help?

OpenStudy (agreene):

12-(-3*-6) 12-(18) -6

OpenStudy (anonymous):

product means to multiply?

OpenStudy (agreene):

yes

OpenStudy (anonymous):

what does quotient mean? i dont understand this one "the quotient of 14 and the different between 7 and -7"

OpenStudy (agreene):

quotient is division. So, with what you wrote: 14/(7--7)

OpenStudy (agreene):

which will give you 14/(7+7) = 14/14= 1/1 = 1

OpenStudy (anonymous):

"the difference between two-thirds of nine and five" is it... 2/3-(9+5)?

OpenStudy (anonymous):

??